OpenStudy (mayah3):

Help please again: 67 = g – (–36)

OpenStudy (turingtest):

answers or how it's done?

OpenStudy (turingtest):

which do you need?

OpenStudy (mayah3):

Both

OpenStudy (mayah3):

But answer first please :)

OpenStudy (turingtest):

31

OpenStudy (turingtest):

ok a negative times a negative is a positive, right? so 67=g-(-36)=g+(-1)(-36)=g+36 subtract 36 from BOTH sides g=67-36=31

OpenStudy (mayah3):

Ohh okay thanks so much :D

OpenStudy (turingtest):

no prob